OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible (2-x)e^x dx

OpenStudy (anonymous):

Notice that our \(2-x\), if differentiated, will reduce to a constant; our \(e^x\), however, will not. This means we should should treat our integral as (by integration by parts): \(\int(2-x)\,\mathrm{d}(e^x)=(2-x)e^x-\int e^xd(2-x)\)

OpenStudy (anonymous):

so I dont need to worry about the (2-x)?

OpenStudy (anonymous):

\[\int (2-x)e^x dx=2\int e^xdx -\int xe^xdx\] is a start

OpenStudy (anonymous):

first part is pretty obviously \(2e^x\) it is the second integral \[\int xe^xdx\] that requires parts

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

is it uv-vdu?